### v1

# Parallel Batch-Dynamic Maximal Matching with Constant Work per Update

Abstract.
We present a work optimal algorithm for parallel fully batch-dynamic maximal matching against an oblivious adversary. In particular it processes batches of updates (either insertions or deletions of edges) in constant expected amortized work per edge update, and in $O(\log^{3}m)$ depth per batch whp, where $m$ is the maximum number of edges in the graph over time. This greatly improves on the recent result by Ghaffari and Trygub (2024) that requires $O(\log^{9}m)$ amortized work per update and $O(\log^{4}m)$ depth per batch, both whp.
The algorithm can also be used for hyperedge maximal matching. For hypergraphs with rank $r$ (maximum cardinality of any edge) the algorithm supports batches of insertions and deletions with $O(r^{3})$ expected amortized work per edge update, and $O(\log^{3}m)$ depth per batch whp. This is a factor of $O(r)$ work off of the best sequential algorithm, Assadi and Solomon (2021), which uses $O(r^{2})$ work per update. Ghaffari and Trygub’s parallel batch-dynamic algorithm on hypergraphs requires $O(r^{8}\log^{9}m)$ amortized work per edge update whp. We leverage ideas from the prior algorithms but introduce substantial new ideas. Furthermore, our algorithm is relatively simple, perhaps even simpler than the sequential hyperedge algorithm.
We also present the first work-efficient algorithm for maximal matching on hypergraphs. For a hypergraph with total cardinality $m^{\prime}$ (i.e., sum over the cardinality of each edge), the algorithm runs in $O(m^{\prime})$ work in expectation and $O(\log^{2}m)$ depth whp. The algorithm also has some properties that allow us to use it as a subroutine in the dynamic algorithm to select random edges in the graph to add to the matching.
parallel graph algorithms, batch-dynamic algorithms
ccs: Theory of computation Dynamic graph algorithms
1. Introduction
We consider the problem of maintaining the maximal matching of a graph under batch insertions and deletions of edges (fully dynamic), or hyperedges in a hypergraph. There have been several important results on sequential fully dynamic maximal matching in the past fifteen years. Onak and Rubinfeld described an dynamic algorithm with $O(\log^{2}n)$ expected amortized time per update that maintains a constant-approximate maximum matching (Onak and Rubinfeld, 2010). Baswana, Gupta and Sen (BGS) presented the first maximal matching algorithm with polylogarithmic time per update (Baswana et al., 2011). Their algorithm supports single updates (insertions or deletions) in $O(\log n)$ amortized time with high probability (whp). [^1] We say that $f(n)=O(g(n))$ with high probability in $n$ if for all constants $c>1$, there exist constants $n_{0},c_{0}>1$ such that for $n\geq n_{0}$, $Pr[f(n)\leq c_{0}g(n)]\geq 1-n^{-c}$. Solomon (Solomon, 2016) then improved the result to $O(1)$ time per update, again amortized and whp. These results are for standard graphs. Assadi and Solomon (AS) then extended the results to hypergraphs (Assadi and Solomon, 2021). In particular, they present an algorithm that, for a hypergraph where each edge has rank at most $r$, supports insertions and deletions in $O(r^{2})$ per update, amortized and whp. All of these results are against an oblivious adversary who knows the algorithm but must select the sequence of updates without knowing any of the random choices by the algorithm. All of these algorithms maintain a leveling structure that try to keep vertices or edges on a level appropriate for their degree.
Algorithms that are robust against an adaptive adversary, a stronger adversary who can respond to random choices by the algorithm, are either much less efficient, give a weaker approximation ratio, or only report the approximate size of the matching. Liu et al.’s batch-dynamic algorithm has update time $O(\alpha+\log^{2}n)$ per element in the batch, where $\alpha$ is the arboricity of the graph  (Liu et al., 2022). Roghani, Saberi, and Wajc (2022) get a better than 2 approximation ratio, but have $O(n^{3/4})$ worst-case update time (Roghani et al., 2022). [^2] The approximation ratio is compared to the maximum matching. Maximal matching gives a 2-approximation. Wajc (2020) achieves constant expected time but for $(2+\epsilon)$-approximation (Wajc, 2020). Bhattacharya et al. (2024) find the size of a matching with approximation ratio better than 2 in polylog update time whp (Bhattacharya et al., 2024). There is no known sequential dynamic maximal matching algorithm that achieves polylog expected updated time against an adaptive adversary. Thus, for the remainder of this work we focus on algorithms that function against an oblivious adversary.
Ghaffari and Trygub (GT) considered the problem in the parallel batch-dynamic setting (Ghaffari and Trygub, 2024). In this setting, instead of inserting or deleting one edge at a time, the user inserts or deletes, on each step, a batch of edges. The size of the batch can vary from step to step. The advantage of considering batches is that an algorithm can then, possibly, apply the updates in parallel, across multiple processors in much better time than applying each one after the other. Parallel batch dynamic updates have been considered for many graph problems (Liu et al., 2022; Acar et al., 2019; Tseng et al., 2019; Acar et al., 2020; Dhulipala et al., 2021; Anderson et al., 2020; Dhulipala et al., 2020; Anderson and Blelloch, 2021). In these results, the efficiency of such algorithms is measured in terms of the total work required to process the batch (number of primitive steps), and the 0pt (also called span or critical path) of dependences in the computation for each batch. One goal in the design of such parallel batch-dynamic algorithms is to achieve asymptotic work that is as good as the best sequential algorithm, or at least close to as good. Algorithms that match the sequential work are referred to as work efficient. A second goal is to minimize the 0pt. Ideally the 0pt should be polylogarithmic in the size of the problem.
GT build on the algorithm of BGS to develop a parallel batch-dynamic data structure with polylogarithmic 0pt per batch update, whp. However, the work is far from work efficient, requiring $O(r^{8}\log^{9}m)$ work per edge update, whp. On a regular graph this reduces to $O(\log^{9}m)$ work per edge update whp. They ask at the end of their paper, with regards to regular graphs: “Can we get a parallel dynamic algorithm for maximal matching with $O(1)$ amortized work per update and poly($\log n$) 0pt for processing any batch of updates?”
In this paper we present a batch-dynamic algorithm that supports updates (insertions or deletions) in $O(1)$ work per edge update for regular graphs and $O(r^{3})$ work for hypergraphs, both amortized and in expectation. The algorithm runs in $O(\log^{3}m)$ 0pt whp. We therefore answer Ghaffari and Trygub’s question in the affirmative (when $r=2$). In particular we have:
Theorem 1.1.
There exists a randomized algorithm that solves the parallel batch-dynamic maximal matching problem on hypergraphs with $O(r^{3})$ amortized and expected work per edge update. Furthermore, processing each batch requires $O(\log^{3}m)$ 0pt with high probability. Here $r$ is the rank of the hypergraph and $m$ is the maximum number of edges to appear at any point in the update sequence. The batch size can vary across steps.
For $r=2$ (standard graphs), this is constant time per update, which matches the best known sequential algorithm (Solomon, 2016). Thus we can conclude the following corollary.
Corollary 1.2.
There exists a parallel randomized batch dynamic graph maximal matching algorithm with constant update time in expectation and polylog 0pt whp, which is work-optimal.
Because the batch dynamic and parallel models are more general than the dynamic and sequential models, our algorithm subsumes the previous algorithms in those spaces. Because maximal matching is lower bounded by $\Theta(m)$, our algorithm is optimal for sequential static maximal matching, parallel static maximal matching, and sequential dynamic maximal matching.
We note that $r$-approximate set-cover for instances where each element belongs to at most $r$ sets reduces to maximal matching in a hypergraph with rank $r$ (Assadi and Solomon, 2021). In the reduction, sets correspond to vertices in the hypergraph and elements to the hyperedges. Our batch dynamic maximal matching algorithm therefore implies the following.
Corollary 1.3.
There exists a randomized parallel batch-dynamic $r$-approximate set cover algorithm that supports batches of insertions and deletions of elements, each element covered by at most $r$ sets. The algorithm does $O(r^{3})$ work per element update (expected and amortized) and has $O(\log^{3}m)$ 0pt per batch whp, where $m$ is the maximum number of elements to ever appear.
This is the current state of the art for the problem.
Within our algorithm, we use static hypergraph maximal matching as a subprocedure, and our work bound is proportional to the cost of static matching. However, directly translating the existing static parallel matching algorithms (Luby, 1986; Birn et al., 2013; Blelloch et al., 2012) to hypergraphs gives $O(mr^{2})$ work, or $O(mr\log m)$ work (Ghaffari and Trygub, 2024). Thus, to improve our work bounds, we develop the first work-efficient, and work optimal, parallel algorithm for maximal hypergraph matching. This builds on the work of Blelloch, Fineman and Shun (BFS) (Blelloch et al., 2012) and improved bounds by Fischer and Noever (FN)  (Fischer and Noever, 2019). FN show that the dependence 0pt of the simple greedy sequential algorithm for maximal matching is shallow when applied to a random permutation. This also applies to hypergraph matching. We extend the approach of BFS to develop a work-efficient algorithm. The cardinality of an edge in a hypergraph is the number of vertices it is incident on, and the total cardinality of a hypergraph is the sum of the cardinality across the edges.
Lemma 1.4.
There exists a parallel algorithm for finding a maximal matching of a hypergraph with $m$ edges and total cardinality $m^{\prime}$ with $O(m^{\prime})$ expected work and $O(\log^{2}m)$ 0pt whp.
Interestingly, this static maximal matching algorithm has additional important properties we make use of. In particular, because of the random ordering, each edge is selected to be a match at random among its remaining incident edges. For each match, call the remaining incident edges at the time it is matched, including itself, the sample space for that match. We call the size of this sample space the price of the match. We show that for any adversarial (oblivious) sequence of deletions we can pay this price by charging every edge deleted at or before its incident matched edge at most two units of price in expectation. The proof is subtle since the sample spaces themselves depend on the random choices.
1.1. Relationship to Prior Algorithms
Our algorithm borrows many ideas from prior level-based dynamic matching algorithms (Onak and Rubinfeld, 2010; Baswana et al., 2011; Solomon, 2016; Assadi and Solomon, 2021; Ghaffari and Trygub, 2024). It is most similar to the Assadi and Solomon (Assadi and Solomon, 2021) hypergraph algorithm. Like (Assadi and Solomon, 2021) and (Solomon, 2016), and unlike the others it is lazy: although a level accurately represents the original sample size when settled, the degree of an edge can grow much larger before it is resettled. This is what allows it to be work efficient.
All of these prior algorithms have the following pattern. First, observe that inserting edges and deleting unmatched edges is easy; the interesting case is deleting matched edges. By choosing which edges are matched from a large sample space, one can protect them from the oblivious adversary—i.e., one can amortize the expensive cost of deleting the matched edge with the many cheap deletes of unmatched edges in the sample space. To keep track of the sample space associated with a matched vertex, a leveling scheme is kept: higher level means a larger sample space. Starting with the analysis by BGS  (Baswana et al., 2011), all the algorithms distinguish matched edges that are deleted by the user (natural deletions) from those deleted indirectly by the algorithm itself (induced deletions). The induced deletions need to be charged to the natural ones.
Our algorithm differs from prior algorithms in important ways, however, and adds several new ideas to achieve parallel work efficiency. One of the most important differences is the random sampling approach. Instead of selecting a single edge among the neighbors of a vertex (or in some special cases in (Assadi and Solomon, 2021), an edge), it selects the samples using a random greedy maximal matching algorithm, as described above. We show that the randomness in this very natural algorithm (the trivial greedy maximal matching on a random ordering of the edges) is sufficient.
An important difference in the analysis of our algorithm is how we charge induced deletions to natural deletions. In the prior algorithms it is done by charging to each newly matched edge some number of deleted edges on a lower level. Ultimately these are charged to a natural deletion. They therefore carefully have to ensure that a matched edge only induces deletes on strictly lower levels. We instead allow a newly matched edge at a low level to induce a deleted edge at an arbitrarily higher level. Thus, we only make an aggregate argument that upper bounds the total size of the sample space of induced deletions relative to the sample space of newly created matches. Ultimately the total sample space of induced deletions is charged to the sample space of natural deletions.
This leads to another notable change relative to prior hypergraph algorithms (Assadi and Solomon, 2021; Ghaffari and Trygub, 2024). In particular, in their leveling schemes the levels differ by a factor of $\Theta(r)$. In our scheme the levels differ by a factor of $2$. This is important in our charging scheme, as discussed in the analysis. Another difference from the the lazy schemes (Solomon, 2016; Assadi and Solomon, 2021) is that we terminate after at most $O(\log n)$ “recursive levels.” This is important for bounding the depth. In Solomon’s (2016) algorithm (Solomon, 2016), in certain cases the recursion depth is $\Omega(n)$. It is not clear whether Solomon’s charging scheme could handle terminating after $O(\log n)$ rounds, but it is relatively easy to show that our charging scheme does support early termination.
We also differ significantly from Ghaffari and Trygub’s parallel batch-dynamic algorithm (Ghaffari and Trygub, 2024). Their algorithm is based on BGS instead of Solomon’s lazy approach. It is therefore unlikely that GT’s algorithm can be improved to $O(1)$ work per edge for standard graphs, even with improvements to all the components. The GT algorithm also uses a significantly more complicated method for randomly selecting edges. In GT, edges are processed by level top-down, whereas our maximal matching processes all levels together. Furthermore, GT has a bespoke method for finding the matching on each level that involves a triply nested loop. To ensure that not too many edges are selected at the same time in parallel, the edges are sampled at varying density across the iterations of the loops. The analysis is quite involved. Together this leads to the high cost per edge update.
2. Preliminaries
Hypergraph matching
A hypergraph $H=(V,E)$ consists of a set of vertices $V$ and a set of edges $E\subset 2^{V}$. Let $n=|V|,m=|E|$ and $m^{\prime}=\sum_{e\in E}|e|$. The rank of a hypergraph is the maximum cardinality of any of its edges. We will be considering hypergraphs of bounded rank, and will use $r(H)$ or just $r$ to indicate the rank. We say two edges are incident on each other if they share a vertex, and an edge is incident on a vertex if it includes the vertex in its set. On a hypergraph $H=(V,E)$ a set of edges $M\subset E$ is said to be a matching if no two edges in $M$ are incident on each other. It is a maximal matching if all edges in $E\setminus M$ are incident on $M$. We say that an edge or vertex is free if it is not incident on any matched edge.
Parallel Model
We assume the fork-join parallel model with work-0pt analysis (Blelloch, 1996; Blelloch et al., 2020). In the model, a thread can fork off any number of child threads which run in parallel, but asynchronously. When all finish, control returns to the parent, which is the join point. Such forking can be nested. Beyond the fork instruction, each thread executes a standard RAM instruction set. We say there is a dependence between two instructions if they are in the same thread, or one is in an ancestor thread of the other, and otherwise they are concurrent. The cost is measured in terms of work and 0pt. The work is the total number of instructions executed by all threads, and the 0pt (also called span or critical path length) is the longest chain of dependent instructions. We allow for concurrent writes, but the ordering is non-deterministic. In particular, we assume that the values that are read are consistent with some total order of the instructions that is compatible with the partial order defined by the dependences.
A PRAM algorithm with $p$ processors and $t$ time can be simulated on the fork-join model in $O(pt)$ work and $O(t)$ 0pt. A fork-join algorithm with $w$ work and $t$ time can be simulated on the PRAM in $O(w/p+t\log^{*}w)$ time randomly w.h.p, or $O(w/p+t\log\log p)$ time deterministically (Blelloch et al., 2020). Because mapping from the fork-join model to the PRAM only adds at worst a $O(\log^{*}w)$ factor to the time in the randomized case, our results are quite robust across the models.
Dynamic model
We are interested in batch dynamic maximal matching. In particular we support inserting and deleting sets (batches) of edges. Each batch can be of arbitrary size. We also assume we can report for any vertex its matched edge (or none if the vertex is free) in constant expected time, and that we can return the set of matched edges in expected work proportional to the size of the set. We assume edges have unique identifiers so they can be hashed or compared for equality in constant time (even though they might have $r$ endpoints). This is not an unreasonable assumption since one can always keep a hash table that hashes an edge to its label external to our algorithm and then use the labels internally. As with much prior work on dynamic maximal matching (Baswana et al., 2011; Solomon, 2016; Assadi and Solomon, 2021; Ghaffari and Trygub, 2024) our algorithms are randomized and we assume an oblivious adversarial user. The oblivious adversary does not know anything about the random choices made by the algorithm.
Standard Algorithms
We make use of several standard parallel algorithms, which we outline here. We make use of prefix sums and filtering, which both require $O(n)$ work and $O(\log n)$ 0pt (Blelloch, 1993). The filter operation takes an array with a subset of elements marked and returns an array with just the marked elements, maintaining order among them. We will need to generate a random permutation. For a sequence of length $n$ this can be done in $O(n)$ expected work and $O(\log n)$ 0pt (Gil et al., 1991).
[…]
We also make use of dictionary data structures based on hashing. A parallel dictionary structure supports batches of insertions, deletions, and membership queries. A batch of any of these operations $k$ can be implemented in $O(k)$ expected work and $O(\log^{*}k)$ 0pt with whp (in $k$(Gil et al., 1991). This assumes the dictionary is large enough to hold the elements (i.e., the load factor is constant). We will need growable/shrinkable dictionaries, but given that we are happy with amortized bounds this can be implemented using the standard doubling/halving tricks (i.e., when too full, double the size and copy the old contents into the new, and when too empty, half the size and copy). In our analysis we will assume that batch of $k$ operations can be applied to a dictionary of size $n$ in $O(k)$ expected amortized work and $O(\log(n+k))$ 0pt whp. Also beyond membership queries we will need to extract a sequence of all the current elements. This can be performed with $O(n)$ expected work and $O(\log(n+k))$ 0pt whp. The size of each dictionary is $O(n)$.
Finally we will need a findNext function that, given an index $i$ in an array, finds the next index $j$ that satisfies some predicate. It needs to run in $O(j-i)$ work and $O(\log(j-i))$ 0pt. This can be implemented using doubling then binary search. On each round $k$, search the next $2^{k}$ elements. In a round, in parallel each element in the range checks the predicate and writes to some flag if satisfied. If none is found in a round, then go to the next round. Each round takes $O(2^{k})$ work and $O(1)$ 0pt. When a round succeeds, we then apply binary search over that range. In each step of binary search, we check if the left half of the list contains an alive element by concurrently writing to a flag. If the left half contains an alive element, we recurse on the left half; otherwise, we recurse on the right half. Thus the index of the next element that satisfies the predicate can be identified in $O(2^{k})$ work and $O(k)$ 0pt. The total work and 0pt is therefore within the required bounds.
Parallel insertions, deletions and increments
In our pseudocode, for simplicity, we often insert and delete elements from various sets in parallel loops. In our parallel model, this would require first applying a groupBy to gather elements based on the set they are updating, and then applying a batch update to each such set. Similarly, in the updateTop code in Section 3, we increment various counters in parallel. This can be implemented with a sumBy. Both of these require at most expected amortized work proportional to the number of updates, and logarithmic 0pt whp. The amortization comes from possibly needing to grow or shrink some of the hash tables for the sets.
3. Random Maximal Hyperedge Matching
In our dynamic algorithm, we must find a maximal matching on a subgraph in two situations: when inserting edges and during random settles. When inserting edges, any fast maximal matching would suffice. However, for the random settles, we need, among other properties, for the matched edges to be picked with sufficient randomness to protect them from the oblivious adversary. For brevity we use the same greedy algorithm for both cases. Sequential (static) greedy maximal matching randomly permutes the edges of a graph, then passes through them once, selecting an edge for the matching if all its endpoints are free (see pseudocode below). The initial permutation, which is chosen uniformly at random, can be generated in linear work and log 0pt (Gil et al., 1991). Parallel greedy maximal matching returns the same matching as the sequential greedy algorithm but processes many edges at once. Consider the dependence graph on the edges, where an edge $e$ depends on $e^{\prime}$ if $e$ is incident to $e^{\prime}$ and has lower priority than $e^{\prime}$. We call an edge a root of the dependence graph if it depends on no edge. In each round of the parallel algorithm, we add all roots to the matching, remove the matched edges and their neighbors, and update the root set. BFS provides a work efficient algorithm (for $r=2$) for matching in this way (Blelloch et al., 2012). In BFS they show that the dependence graph has 0pt $O(\log^{2}n)$ (Blelloch et al., 2012); this was improved to $O(\log n)$ by FN (Fischer and Noever, 2019).
However, a direct translation of BFS (Blelloch et al., 2012) gives $O(mr^{2})$ work for hypergraphs. We modify the method by which the root set is updated to yield $O(m^{\prime})$ work, where $m^{\prime}$ is the total cardinality of the hypergraph. We keep track of the following data structures. For each vertex $v$, neighbors$(v)$ gives a linked list of incident edges, edges$(v)$ gives an array of edges sorted by priority, and top$(v)$ gives the index of the highest priority remaining edge in edges$(v)$. We do not remove edges from edges$(v)$; instead, we increase top$(v)$ to indicate that prior edges are deleted. Each edge $e$ has pointers to each of its instances (up to $r$ each) in neighbors$(v)$ and edges$(v)$, and an atomic counter (counter), initially set to 0, indicating on how many vertices it is the highest priority remaining edge. We set done$(e)$ to be true if $e$ is being deleted this round, and false otherwise. Note that an edge is a root iff it is the highest priority remaining edge with respect to all of its vertices. When edges(v)[top$(v)$] is set to be deleted, we call updateTop$(v)$, which finds $v$’s new highest remaining edge $e^{\prime}$ and increments counter$(e^{\prime})$. If counter$(e^{\prime})$ equals the rank of $e^{\prime}$, we add it to the new root set. We repeat on the new root set and remaining graph. The overall algorithm is written in the function parallelGreedyMatch.
--Input: A Graph $G=(V,E)$
--Output: A matching augmented with sample spaces
sequentialGreedyMatch$(G=(V,E))$:
$\pi\leftarrow$ getRandomPermutation()
$E^{\prime}\leftarrow E\mbox{ radix sorted by }\pi$
$X\leftarrow\emptyset$
for $e\in E^{\prime}$:
if $\mbox{free}(e)$:
$\mbox{free}(e)\leftarrow false$, $S_{e}\leftarrow\{e\}$
for edges $\mbox{edges }e^{\prime}\mbox{ incident to }e$ :
if $\mbox{free}(e^{\prime})$ :
$\mbox{free}(e^{\prime})\leftarrow false$, $S_{e}\leftarrow S_{e}\cup\{e^{\prime}\}$
$X\leftarrow X\cup\{(e,S_{e})\}$
return $X$
--Input: A vertex $v$
--Effect: Corrects top$(v)$ if needed
--Output: Returns edge to add to frontier in needed
updateTop(v) :
if $\mbox{done}(\mbox{edges}(v)[\mbox{top}(v)])$ :
$\mbox{top}(v)\leftarrow\mbox{findNext}(\mbox{top}(v),\mbox{edges}(v))$
if $\mbox{top}(v)\neq|\mbox{edges}(v)|$ :
$e_{t}=\mbox{edges}(v)[\mbox{top}(v)]$
$\mbox{increment}(\mbox{counter}(e_{t}))$
if $\mbox{counter}(e_{t})=|V(e_{t})|$ : return $e_{t}$
return $\bot$
--Input: A Graph $G=(V,E)$
--Effect: Clears $G$
--Output: A matching augmented with sample spaces
parallelGreedyMatch($G=(V,E)$):
$\pi\leftarrow$ getRandomPermutation()
$W\leftarrow\{e\in E\;|\;\forall_{e^{\prime}\in N_{G}(e)}\pi(e)<\pi(e^{\prime})\}$
parfor $v\in V$ :
$\mbox{edges}(v)\leftarrow$ sort $\{e\;|\;v\in e\}\mbox{ by }\pi(e)$
$\mbox{top}(v)\leftarrow 0$
$\mbox{increment}(\mbox{counter}(\mbox{edges}(v)[0]))$
$X\leftarrow\emptyset$
while (|W| > 0) :
$D^{\prime}\leftarrow\mbox{groupBy}(\{(n,w):n\in N_{G}(w)\cup\{w\},w\in W\})$
$\mbox{minIn}\leftarrow\{(\arg\min_{e\in F}\pi(e)),n):(n,F)\in D^{\prime}\})$
$X\leftarrow X\cup\mbox{groupBy}(\mbox{minIn})$
$\mbox{finished}\leftarrow W\cup N_{G}(W)$
parfor $e\in\mbox{finished}$ : $\mbox{done}(e)\leftarrow true$
$V_{f}\leftarrow\bigcup_{e\in\mbox{finished}}\{v|v\in e\}$
$W\leftarrow\{\mbox{updateTop}(v):v\in V_{f},\mbox{updateTop}(v)\neq\bot\}$
$E\leftarrow E\setminus\mbox{finished}$
return $X$
Figure 1. Sequential and parallel greedy maximal matching.
Define the sample space $S_{e}$ of a matched edge $e$ to be a set of unmatched edges we are assigning to $e$ [^3] The sample space corresponds to the temporarily deleted edges from Assadi and Solomon’s algorithm (Assadi and Solomon, 2021). The maximal matching we return with parallelGreedyMatch is augmented with sample spaces. For our dynamic algorithm to be efficient, we need for the sample spaces to have certain properties. To cover the cost of matching, we need all edges to be assigned to a sample space. We need an edge to belong to at most one sample space, so that its deletion only counts against the life of one epoch. Thus, we need the sample spaces to be disjoint. To ensure that edges are not reused in other sample spaces, we implicitly remove them from the graph until the matched edge is deleted. Thus we need for sampled edges to intersect their chosen matched edge, so that removing them does not violate maximality in future time steps. We formalize these desiderata in the following lemma.
Lemma 3.1.
The function parallelGreedyMatch has the following properties.
  1. (1)
    $\{S_{e}:(e,S_{e})\in X\}$ partitions $E$.
  2. (2)
    For $(e,S_{e})\in X$, for all $e^{\prime}\in S_{e}$, $e\cap e^{\prime}\neq\emptyset$.
  3. (3)
    $\{e:(e,S_{e})\in X\}$ is a maximal matching on $G$.
Proof.
We can analyze sequentialGreedyMatch instead of parallelGreedyMatch because, as shown by BFS, they process edges in the same order (Blelloch et al., 2012).
  1. (1)
    Note that if $e^{\prime}$ is added to $S_{e}$ in time step $t$, $\mbox{free}(e^{\prime})=true$ before time step $t$ and $\mbox{free}(e^{\prime})=\mbox{false}$ after time step $t$. Thus the $S_{e}$ are disjoint.
  2. (2)
    Observe that if $\mbox{free}(e^{\prime})=\mbox{false}$, then $e^{\prime}\in S_{e}$ for some $e$. Consider the state of $\mbox{free}(e^{\prime})$ at time step $t$. If $\mbox{free}(e^{\prime})$ is true, $e^{\prime}$ is added to the matching, so $e^{\prime}\in S_{e^{\prime}}$. Otherwise, $e^{\prime}$ is already in some sample space $S_{e}$. Thus every edge belongs to a sample space. Thus the sample spaces partition $E$.
  3. (3)
    Note that we only add edges to the sample space that are incident on the matched edge.
Lemma 3.2.
All of the calls to updateTop cost, in aggregate, $O(m^{\prime})$ work, where $m^{\prime}$ is the total cardinality of the edges. Each call individually is $O(\log m)$ 0pt.
Proof.
Each call to updateTop$(v)$ slides $v$’s pointer along its edge list to the right, to the new highest priority remaining edge. Using the findNext operation, a call that slides a pointer $d$ distance costs $O(d)$. Because there are $m^{\prime}$ total edges across the lists, the pointers slide a total of $O(m^{\prime})$ distance, for $O(m^{\prime})$ work. Since the edge lists have at most $m$ members, each findNext call takes $O(\log m)$ 0pt, so updateTop has $O(\log m)$ 0pt. ∎
Theorem 3.3.
parallelGreedyMatch takes $O(m^{\prime})$ expected work and $O(\log^{2}m)$ 0pt whp.
Proof.
Because the edges are given random priorities, bucket sorting takes $O(m)$ work in expectation and $O(\log m)$ 0pt whp (Blelloch et al., 2012; Cormen et al., 2001). Each edge $e$ is deleted once and costs $|e|$ work outside of updateTop, for $O(m^{\prime})$ work. In aggregate, internal calls to groupBy cost no more than $O(m)$ in expectation. Since updateTop in aggregate costs $O(m^{\prime})$, we have $O(m^{\prime})$ work total. Since updateTop and filter have $O(\log m)$ 0pt, each round of GreedyMatch is $O(\log m)$ 0pt. By Fischer and Noever there are $O(\log m)$ rounds whp (Fischer and Noever, 2019), for $O(\log^{2}m)$ 0pt total. ∎
3.1. Random selection in Greedy Maximal Matching
In Solomon’s algorithm (Solomon, 2016), about half of the sampled edges get deleted before the matched edge, because the matched edge was chosen uniformly at random from the sample set. This property is crucial for the analysis, because it shows that an expensive matched edge deletion can be amortized against many cheap updates (the unmatched deletions before it). In Solomon’s algorithm, the sample space is chosen deterministically (based on the present graph state), then a random mate is chosen.
This is not true of parallelGreedyMatch, which chooses a matched edge and its sample space together, in the same random process. Note that we do not tell the adversary the sample space partitioning of a graph. But if we did tell the adversary the sample space of a matched edge, because the adversary knows the structure of the graph, the adversary would learn something about which edge was likely matched. For example, consider the path on 4 vertices $(1,2),(2,3),(3,4)$. Suppose that parallelGreedyMatch is run and the sample space of one of the matched edges is $\{(1,2),(2,3),(3,4)\}$. Since a matched edge is incident on all its sample space it follows that $(2,3)$ was matched. Thus, even knowing the size of the sample space reveals information about which edge was likely matched. We still can demonstrate that parallelGreedyMatch has the necessary randomness to protect its matched edges from the adversary, but we must approach the analysis differently.
To each matched edge $e\in E$, we assign price $|S_{e}|$; unmatched edges are given price 0. Suppose that the user deletes one edge $d_{t}$ per time step $i$ until all of the edges that participate in parallelGreedyMatch are deleted. If $p(d_{t})$, the matched edge whose sample space $d_{t}$ belongs to, has not yet been deleted (or if $d_{t}=p(d_{t})$), we call the delete “early.” Otherwise, we call the user delete “late.” When the user deletes an unmatched edge $e^{\prime}$, we pay 1 and subtract 1 from $p(e^{\prime})$’s price [^4] We only reduce $p(e^{\prime})$’s price if $p(e^{\prime})$ is still present in the graph (i.e., if $e^{\prime}$ is an early delete).. When the user visits a matched edge $e$, we pay the price of $e$. Let $\Phi(e)$ be the total price we pay when visiting edge $e$.
Remark: Consider a user delete of edge $d_{t}$. Suppose that $d_{t}$ is a late delete. Since $d_{t}$’s matched edge is gone, $d_{t}$ is unmatched. Thus while visiting $d_{t}$ we will pay 1 unit of price. Thus $\mathbb{E}[\Phi(d_{t})]=1$. Note that we do not need this fact, as late deletes do not appear in our analysis later.
Lemma 3.4.
If $d_{t}$ is an early delete, then $\mathbb{E}[\Phi(d_{t})]\leq 2$.
Proof.
We consider sequentialGreedyMatch because it processes edges in the same order as the parallel version (Blelloch et al., 2012). Let UserRemain be the remaining incident edges on $d_{t}$ at (the beginning of) time $t$. The other edges that were incident on $d_{t}$ at time 0 have been deleted by the user. Let $\Pi$ be the set of permutations (that we would input into sequentialGreedyMatch) such that $d_{t}$ is an early delete. Then $\Pi$ is the set of permutations such that $p(d_{t})\in\mbox{UserRemain}$. Let $r$ be the time step in sequentialGreedyMatch at which $d_{t}$ is marked to be not free (the time step where $d_{t}$ is effectively removed from the graph). [^5] Note that we have two distinct uses of time: the time in the user delete sequence, and the time in the run of sequentialGreedyMatch. Note that the sequentialGreedyMatch time happens completely before the user delete sequence time. Note that these times are separate from each other: if $t$ is large, it is possible that $r$ is large or small. Let PermRemain be the set of edges incident on $d_{t}$ at step $r$. Group $\Pi$ into equivalence classes based on, for $\pi\in\Pi$, the value of $[r,(\pi_{i})_{i=1}^{r-1}]$. Note that PermRemain is determined by $[r,(\pi_{i})_{i=1}^{r-1}]$. Let $\pi\in\Pi$, and let $B$ be the equivalence class for which $\pi\in B$.
We consider the possible values of $\pi^{-1}(r)$. Since $d_{t}$ must delete in step $r$, $\pi^{-1}(r)$ must be in PermRemain. Note that $\pi^{-1}(r)=p(d_{t})\in\mbox{UserRemain}$. Since there are no other restrictions, we have that $\pi^{-1}(r)\in\mbox{PermRemain}\cap\mbox{UserRemain}$. Thus $|B|=|\mbox{PermRemain}\cap\mbox{UserRemain}|(m-r)!$. Next, we consider the value of $\Phi(d_{t})$ if $d_{t}$ is matched. At user time 0, $d_{t}$ had price —PermRemain—. However, each unmatched edge in $d_{t}$’s sample space that the user visits reduces $d_{t}$’s price by 1. Thus, the price of $d_{t}$ is $|\mbox{PermRemain}|-|\mbox{PermRemain}\cap\overline{\mbox{UserRemain}}|=|\mbox{PermRemain}\cap\mbox{UserRemain}|$. There are $(m-r)!$ permutations in $B$ where $d_{t}$ matches at time step $r$.
Drawing a permutation uniformly at random from $B$, we get $\mathbb{E}[\Phi(d_{t})]=|\mbox{PermRemain}\cap\mbox{UserRemain}|(1/|\mbox{PermRemain}\cap\mbox{UserRemain}|)+(1-1/|\mbox{PermRemain}\cap\mbox{UserRemain}|)1\leq 2$. Since $B$ was an arbitrary equivalence class we conclude that $\mathbb{E}[\phi(d_{t})]=\mathbb{E}[\phi(d_{t})|\pi\in B]\leq 2$. ∎
Note that Lemma 3.4 is true for all times $t$; this means that, regardless of the deletes the user chooses, the next delete has constant expected price. Even though the price of later deletes depends on the price of earlier deletes, the expectation remains no more than 2. We also note that, when a graph is fully deleted, the sum of the prices of early deletes is the number of edges in the graph. Let $\Phi^{\prime}(d_{t})=\begin{cases}\Phi(d_{t})&d_{t}\textrm{ is early}\\ 0&\textrm{otherwise}\end{cases}$.
Lemma 3.5.
The early deletes on the sample space of deleted matched edge $e$ contribute $|S_{e}|$ total price. Furthermore, $\sum_{i=1}^{m}\Phi^{\prime}(d_{t})=m$.
Proof.
Observe that if a matched edge $e$ has price $|S_{e}|-j$ on deletion, then $j$ unmatched edges in $S_{e}$ were deleted, and the updates deleting these $j$ unmatched edges paid $j$ price. Thus the early deletes on the sample space of $e$ contribute $|S_{e}|$ total price. Thus the price of all early deletes is $m$. ∎
isHeavy$(e)$ : return $|C(e)|\geq 4r^{2}2^{l(e)}$
--Input: an edge $m$ along with a set of incident edges $S_{e}$
--Effect: $m$ is added as a match with sample edges $S_{e}$
addMatch$(m,S_{e})$ :
$\mbox{insert}(M,m)$
parfor $e\in S_{e}$ :
type$(e)\leftarrow$sampled
$p(e)\leftarrow m$
type$(m)\leftarrow$matched
parfor $v\in V(m)$ : $p(v)\leftarrow m$
$S(m)\leftarrow S_{e}$
$C(m)\leftarrow\emptyset$
--Input: a matched edge $m$
--Assumes: sample edges are converted to cross edges
--Effect: removes $m$’s owned edges and frees vertices
--Output: the cross edges owned by $m$
removeMatch$(m)$ :
$\mbox{delete}(M,m)$
$E\leftarrow C(m)$
parfor $v\in V(m)$ : $p(v)\leftarrow\bot$
parfor $e\in E$ : removeCrossEdge$(e)$
return $E$
--Input: an edge $e$ that is not in the structure
--Effect: add $e$ to the structure as a cross edge
addCrossEdge$(e)$ :
type$(e)\leftarrow$cross
$v_{max}\leftarrow argmax_{v\in V(e)}l(p(e))$
$e^{\prime}\leftarrow p(v_{max})$
insert$(C(e^{\prime}),e)$
parfor $v\in V(e)$ : insert$(P(v,l(e^{\prime})),e)$
--Input: a cross edge $e$
--Effect: removes $e$ from the structure
removeCrossEdge$(e)$ :
$C(p(e))\leftarrow C(p(e))\setminus\{e\}$
parfor $v\in V(e)$ : delete$(P(v,l(p(e))),e)$
type$(e)\leftarrow$unsettled
--Input: a set $E$ of newly matched edges
--Effect: adjusts cross edge incident on $E$ to their new level
adjustCrossEdges$(E)$ :
$V\leftarrow\bigcup_{e\in E}V(e)$
$C\leftarrow\bigcup_{v\in V}\bigcup_{i\in[0,l(p(v)))}P(v,i)$
parfor $e\in C$ : removeCrossEdge$(e)$
parfor $e\in C$ : addCrossEdge$(e)$
--Input: the cross edges owned by removed heavy matches
--Effect: remove and rematch possibly breaking matches
--Output: the cross edges owned by new removed matches
randomSettle$(E)$ :
$G^{\prime}\leftarrow(V(E),E)$
$X\leftarrow$parallelGreedyMatch$(G^{\prime})$
stolen $\leftarrow\{p(v):(m,\_)\in X\;|\;\exists_{v\in V(m)}(p(v)\neq\bot)\}$
parfor $(e,S_{e})\in X$ : addMatch$(e,S_{e})$
adjustCrossEdges($(\{m:(m,\_)\in X\})$
bloated $\leftarrow\{m:(m,\_)\in X\;|\;\mbox{isHeavy}(m)\}$
[…]
--Input: a set of matched edges
--Effect: remove matches, rematch the light matched edges
--Output: the cross edges owned by the heavy matches
deleteMatchedEdges$(E)$ :
parfor $e\in\bigcup_{m\in E}S(m)$ : addCrossEdge$(e)$
heavy $\leftarrow\{m\in E\;|\;\mbox{isHeavy}(m)\}$
light $\leftarrow E\setminus\mbox{heavy}$
insertEdges$(\bigcup_{m\in\mbox{light}}\mbox{removeMatch}(m))$
return $\bigcup_{m\in\mbox{heavy}}\mbox{removeMatch}(m)$
--User Interface: insert a batch of edges
insertEdges$(E)$ :
free $\leftarrow\{e\in E\;|\;\forall_{v\in e}(p(v)=\bot)\}$
$G^{\prime}\leftarrow(V(free),free)$
[…]
parfor $e\in Y$ : addMatch$(e,\{e\})$
parfor $e\in E\setminus Y$ : addCrossEdge$(e)$
--User Interface: delete a batch of edges
deleteEdges$(E)$ :
matched $\leftarrow\{e\in E\;|\;\mbox{type}(e)=\mbox{matched}\}$
parfor $e\in E\setminus\mbox{matched}$ :
if $\mbox{type}(e)=\mbox{cross}$ : removeCrossEdge$(e)$
else : $\mbox{delete}(S(p(e)),e)$
parfor $m\in\mbox{matched}$ : delete$(S(m),m)$
$E^{\prime}\leftarrow$deleteMatchedEdges(matched)
sampledEdges $\leftarrow 0$
while $(2|E^{\prime}|>\mbox{sampledEdges})$ :
sampledEdges $\leftarrow\mbox{sampledEdges}+|E^{\prime}|$
$E^{\prime}\leftarrow$randomSettle$(E^{\prime})$
insertEdges$(E^{\prime})$
Figure 2. Parallel batch-dynamic algorithm for dynamic matching with $O(r^{3})$ expected amortized cost per edge update.
[…]
$p(e)$ owner of edge $e$
$V(e)$ vertices of edge $e$
$S(m)$ sample edges owned by matched edge $m$
$C(m)$ cross edges owned by matched edge $m$
$l(m)=\lfloor\lg|S(m)|\rfloor$ level of matched edge $m$
$p(v)$ matched edge incident on $v$, or $\bot$ if none
$P(v,l)$$\{e\in\mbox{cross edges}\;|\;p(e)=l\wedge v\in e\}$
Table 1. The notation used in the algorithm.
4. Algorithm
We now describe our parallel batch-dynamic algorithm. Roughly, the algorithm follows the structure of prior level-based dynamic maximal matching algorithms (Baswana et al., 2011; Solomon, 2016; Assadi and Solomon, 2021; Ghaffari and Trygub, 2024). In these algorithms, and ours, inserting edges and deleting unmatched edges is reasonably straightforward, and the difficulty comes in deleting matched edges. When deleting a matched edge we need to find if any of the incident edges can be matched. If the degree of the edge is high this can be costly. The idea of the prior algorithms is to randomly select a possible edge to match across a sufficiently large sample (roughly its degree). The oblivious adversary therefore does not know which of the sampled edges is selected and in expectation will delete half the edges before reaching the one that was selected. When a matched edge is encountered, the algorithms, roughly speaking, need to explore neighboring edges, but this cost is amortized against the prior cheap deletes of unmatched edges.
The level-based algorithms all use some leveling scheme that consists of a logarithmic number of levels and places every matched edge at a level that is appropriate for its sample size (and degree). The algorithms then ensured that the cost of finding matches in incident edges is either proportional (within factors of $r$) to the sample size for that level, or it is possible to re-sample among a larger set and charge the cost to the new sample. Finding an edge to add to a matching is referred to as settling.
Our algorithm most resembles Assadi and Solomon’s hyperedge algorithm (Assadi and Solomon, 2021), although, beyond being parallel, it differs in important ways that are described in Section 1.1.
We now describe how our leveling scheme works and outline the algorithm and analysis. We will describe the analysis in detail in Section 5. Our leveling structure is defined as follows.
Definition 4.1.
A leveled matching structure for an undirected hypergraph $G=(V,E)$ and matching $M\subseteq E$ maintains the following invariants:
  1. (1)
    All edges of $E$ are either cross edges or sampled edges. $M$ is a subset of the sampled edges.
  2. (2)
    Every edge is owned by an incident matched edge (a matched edge owns itself).
  3. (3)
    Every matched edge $e$ that owns $s$ sample edges is assigned a level $l=\lfloor\lg s\rfloor$.
  4. (4)
    The owner $m$ of any cross edge $e$ must be on the maximum level of any matched edges incident on $e$.
The second invariant implies the matching is maximal. Throughout we use $\lg n$ to mean $\log_{2}n$.
The invariants above are maintained between every batch operation. During the operation itself, some can be violated. In particular, updates will involve “unsettling” edges to be resettled.
To efficiently access a leveled matching structure we will use the following data structures.
  • We maintain the set of matched edges $M$.
[…]
  • For every matched edge $m$ we further maintain a set of its sampled edges ($S(m)$) and a set of its owned cross edges ($C(m)$).
  • For every vertex $v$ we maintain $p(v)$ that specifies the matched edge that covers it, if there is one, or $\bot$ if there is none. We also maintain a mapping ($P(v,l)$) from level $l$ to a set of cross edges. The set will contain all cross edges $e$ at level $l=l(e)$ that are incident on $v$.
Table 1 summarizes the data structures and other notation. Figure 2 lists the pseudocode for our algorithm. The user-level functions are insertEdges and deleteEdges, which insert and delete, respectively, batches of edges. Four of of the functions are just for adding and removing matched and cross edges. They just update the relevant data structures to maintain the invariants. When adding a match we also add all its sample edges. When removing a match, we remove its cross edges and mark its vertices as free. Removing a match assumes the sample edges have already been converted to cross edges.
Inserting a batch of edges first identifies which of the edges are free (are not incident on a matched edge) and runs a maximal matching on these. This could be any maximal matching. It then adds the matched edges to the leveled matching structure, each with just themselves as the sample. This places them on level 00 since $\lg 1=0$. Any remaining edges are added as cross edges.
Deleting a batch of edges has to separate the edges that are matched from those that are not. For those that are not we need to update the data structures to remove them. These unmatched edges could either be sample edges or cross edges, but both types are easy to remove. The bulk of the algorithm is for processing the deleted matched edges. To handle matched edges we first remove them from their sample space because they no longer exist in the graph and should not be reinserted. Then we call deleteMatchedEdges, which is the workhorse of the algorithm.
The function deleteMatchedEdges first takes all the sample edges owned by the matched edges to be removed and inserts them as cross edges. We note that when inserting them as cross edges they might or might not be owned by the same matched edge. The algorithm then separates the matched edges into two kinds: light and heavy. The light matches are ones for which the number of cross edges it owns is not much larger than its sample size, and in particular for a match $e$, less than $4r^{2}2^{l(e)}$. We allow ourselves to charge work proportional to the sample size to each match, and hence can afford to rematch these owned edges directly. The algorithm therefore simply removes the light matched edges using removeMatch, which also removes all their owned cross edges. The algorithm then reinserts the cross edges with insertEdges. [^6] In Solomon’s algorithm this is referred to as deterministic settling. The insertEdges will find new matches for the removed cross edges, if possible.
We are now left with deleting the heavy matches. We cannot afford to find matches for the cross edges owned by the heavy matched edges without taking new samples, which is performed by random settling. Before random settling, we remove from the structure all the heavy matches along with their owned cross edges (using removeMatch). Random settling is then performed in rounds. One round, i.e., one call to randomSettle, first takes all cross edges owned by the heavy matches and runs a random greedy matching on them. The output of this matching will be a maximal set of matched edges, along with a sample set for each. Each matched edges $e$ is added to the data structure with their sample $S_{e}$ at the appropriate level ($\lfloor\lg|S_{e}|\rfloor$) using addMatch.
At this point there are a couple of invariants the random settle has broken that need to be fixed. Firstly, the levels of matches have changed so some of the cross edges incident on the new matches may no longer satisfy Invariant 4.1.4—i.e. they might be owned by a match that is not at the maximum level of their incident matched edges. Secondly, since we selected new matches among all edges owned by a match (not just the free ones), the new matches might be incident on existing matches. This would violate the matching property. We call these existing matches stolen matches since the algorithm will need to delete them.
Random settle fixes the cross edges by calling adjustCrossEdges. This removes all cross edges that might change and then puts them back in with ownership on the maximum incident level. This makes use of the $P(v,l)$ data structure. In particular, for all the vertices that belong to the new matches it collects all the cross edges up to one less than the level of the new match. These are cross edges that must be owned by the new match to satisfy Invariant 4.1.4. The cost to transferring ownership can be charged to the new match the ownership is assigned to. If this cost charged to a match is significantly larger than the sample size of the match, however, the match cannot afford the charge. In particular this is the case if the match $m$ now owns at least than $4r^{2}2^{l(e)}$. We refer to these as the bloated matches, and these will be deleted and their owned edges random settled in the next round.
At the end of a round of random settling we therefore have two sets of matched edges to be deleted: the bloated and the stolen edges. We delete these edges using deleteMatchedEdges. As already described, this will partition them into light and heavy [^7] The bloated edges will always be heavy. The light ones can be directly removed and their owned edges reinserted with insertEdges. The heavy ones are removed and their owned edges are returned to be matched in the next round of random settle.
The rounds of random settling continue until either a round is empty, or two times the number of edges to process in a round is less than the total number of edges processed in prior rounds. In the second case, we insert the remaining edges using insertEdges, and the cost is charged against the prior work. Because of the termination condition, each round will double in size. Since the maximum size of a round is $m$, the total number of edges in the graph, there can be at most $O(\log m)$ rounds.
5. Analysis
We now analyze the cost bounds. We consider both work and 0pt, although the 0pt analysis is straightforward and left until the end. As in prior dynamic maximal matching algorithms (Baswana et al., 2011; Solomon, 2016; Assadi and Solomon, 2021; Ghaffari and Trygub, 2024), it is useful to consider the lifetime of a match from when it is created until it is deleted. This is referred to as an epoch. It is also useful to distinguish epochs that are deleted by the user in the call to deleteEdge, vs. ones that are deleted internally by the algorithm by a random settle. The first kind are referred to as natural deletions of an epoch and the second as induced deletions of an epoch. In the following we will equate the epoch $e$ with the edge $e$ corresponding to its match.
Roughly speaking, the work of the natural deletions of epochs can be covered by the fact that on average about half the sampled edges will be deleted by the user before a matched edge is selected. [^8] We intentionally do not say “in expectation” here, since the argument is more subtle. However, the induced deletions of epochs cannot be charged in this way since these epochs are deleted prematurely, possibly immediately after they are created. We therefore charge induced deletions to the creation of new epochs, and ultimately to a natural deletion. We will therefore consider two work costs to associate with epochs. The first is the direct work which is associated with every epoch $e$ and covers all the work for deleting matched edges. Furthermore all work of each batch operation on deleting matched edges can be associated with epochs created and deleted during that operation. The second is the total work, which is only associated with natural epochs, and charges all the direct work of induced deletions to the natural deletions. We will show that the expected direct work for each epoch $e$ (natural or induced) is $O(r^{3}2^{l(e)})$ and the expected total work for each natural epoch $e$ is $O(r^{3}2^{l(e)})$.
For each step in the algorithm, our charging scheme deterministically charges it to an epoch (or set of epochs). However the amount of work charged to an epoch is in expectation because the subroutines of the algorithm (static maximal matching, groupBy) themselves are randomized and have expected work bounds. Thus, many of the lemmas that follow have work bounds in expectation.
5.1. Direct Work: Charging to all Epochs
We start by analyzing the direct work that we charge to each epoch. Here we only consider the work when the user deletes matched edges. The work for adding edges and deleting unmatched edges is just $O(r)$ per edge and is left to the end.
We partition the direct work into three components: the work associated with the light edges when deleting matched edges (the light work), the work associated with random settling (the heavy work), and the work for the final insertEdges (the final work). The light work is charged to the deletion of epochs while the heavy and final work are charge to the creation of epochs. Specifically, the light work includes: all work in insertEdges when called from deleteMatchedEdges, as well as the work of converting sample edges to cross edges, and all work on light edges in deleteMatchedEdges. The final work just includes the final call to insertEdges in deleteEdges. All other work is heavy work. This includes the work in random settle and the work in deleteMatchedEdges associated with heavy edges. We first analyze insertEdges since it is part of the light work.
Lemma 5.1.
The expected work of insertEdges$(E)$ is $O(r|E|)$.
Proof.
The random greedy match does $O(\sum_{e\in E}|e|)\leq O(r|E|)$ expected work. Determining if a vertex is free, adding a match, and adding a cross edge each require $O(r)$ work, and there are $O(|E|)$ such operations for a total of another $O(r|E|)$ work. ∎
Lemma 5.2.
The light work can be covered by charging $O(r^{3}2^{l(e)})$ in expectation to each deleted match $m\in\textsc{light}$ in deleteMatchedEdges.
Proof.
The work in deleteMatchedEdges for converting the sample edges to cross edges is $O(r)$ per edge. Each level $l$ light edge has at most $O(2^{l(e)})$ edges in its sample, for $O(r2^{l(e)})$ charge to each light edge.
The input edges $E$ to insertEdges when called from deleteMatchedEdges are the union of the owned edges of light edges. Each $e\in\mbox{light}$ contributed at most $O(r^{2}2^{l(e)})$ edges to $|E|$ since it was light. By the previous lemma the expected work of insertEdges is $O(r|E|)$, or $O(r)$ per edge. Since $O(r^{2}2^{l(e)})$ edges each charge $O(r)$ to their deleted light matched edge, a deleted light matched edge in total gets charged $O(r^{3}2^{l(e)})$. ∎
The random settle proceeds in rounds and charging goes across adjacent rounds so we consider the full cost across all rounds.
Lemma 5.3.
The heavy work across all rounds of a batch update can be covered by charging $O(r^{3}2^{l(e)})$ in expectation to each newly matched epoch on an edge $e$ during the batch update.
Proof.
The cost of each round of random settle is either covered by matches in the current round or matches in the next round. Specifically, all but raising of matches edges (in raiseMatches) that end up being bloated, and the cost of deleteMatchedEdges at the end, will be covered by the current round.
The total sample size in a round (with input $E$) is $|E|$ since all edges are in a sample. The random greedy match does $O(r|E|)$ expected work by Theorem 3.3. The cost of adding matches is $O(r)$ per edge in the sample sets ($S_{e}$) and hence is $O(r|E|)$. The work for determining which incident edges need to be deleted is also $O(r|E|)$. The cost of raising matches is proportional to the number of cross edges that are raised. Each matched edge $e$ that does not become bloated will collect at most $O(r^{2}2^{l(e)})$, and each such edge will have cost $O(r)$ for a total of $O(r^{3}2^{l(e)})$ work associated with $e$.
The cost for raising edges that become bloated will be charged to the next round. In particular, if there are $E_{next}$ edges in the next round, at most $|E_{next}|$ cross edges will be raised and become bloated, since each will contribute to the next round. Each such edge contributes $O(r)$ work for $O(r|E_{next}|)$ total. The cost of deleteMatchedEdges is split between light and heavy work, as discussed before. In particular any of the light edges are charged to light work. The cost of processing heavy edges is at most $O(r|E_{next}|)$ in expectation since each such edge and its owned edges are passed to the next round.
We also need to cover the work of the initial call to deleteMatchedEdges in deleteEdges. Any work on light edges is charged to light work, and any work on heavy edges is covered by matching the edges $E$ passed to the first round of random settle.
All charges except for raising matches that are light have cost of $O(r|E_{i}|)$ charged against some round $i$. This is $O(r)$ per edge in the sample sets of new epochs, and hence associates $O(r2^{l(e)})$ with each new epoch on edge $e$. The cost is therefore dominated by the $O(r^{3}2^{l(e)})$ work associated with raising the matched edge $e$ when it ends up light. ∎
Lemma 5.4.
The final work can be covered by charging $O(r2^{l(e)})$ in expectation to each newly matched epoch during a batch update.
Proof.
The total number of edges in the samples up to the final step is sampledEdges. The insertEdges at the end on input $E$ requires $O(r|E|)$ expected work by Lemma 5.1. Since by the termination condition we have that $2|E|<\textsc{SampledEdges}$, the final insert cost can be covered by the prior sampled edges during the rounds of random settle. ∎
Putting the work costs together we get the following bound on work.
Lemma 5.5.
All work in deleteEdges to delete matched edges can be covered by charging $O(r^{3}2^{l(e)})$ in expectation to each start or end of an epoch during the batch.
Proof.
Follows from Lemma 5.2, 5.3, and 5.4, which cover all work. ∎
5.2. Total Work: Charging to Natural Epochs
We have so far been able to charge all work in the algorithm to epochs at a rate of $O(r^{3}2^{l(e)})$ per epoch $e$, whether natural or induced. We now need to charge the work associated with induced deletions of epochs to natural deletions of epochs. This will be based on a counting argument where we will show that the total number of edges that are in the sample of natural deletions is at least a constant fraction as many as the total number in induced deletions. This will allow us to cover the cost of the induced deletions.
There are two sources of induced deletions, both caused by a random settle. Firstly, when we randomly match an edge $e$, there can be edges that are already matched and incident on $e$ (stolen in the algorithm) and need to be deleted. We refer to these as stolen deletes. Then there are the deletes that are due to the fact that after raising an edge it might be too heavy for its level (bloated deletes). In the analysis we cover the stolen deletes in the current round as well as the bloated deletes from the previous round with new matches in the current round.
We say that the deleted sample size of a round is the total sample size of stolen deletes from the current round and bloated deletes from the previous round. We say that the added sample size of a round is the total sample size of newly matched edges in the current round.
Lemma 5.6.
On each round of randomSettle with deleted sample size $S_{d}$ and added sample size $S_{a}$, $S_{a}\geq 2S_{d}$.
Proof.
The input edge set $E$ to randomSettle is the union of the owned cross edges of a set of heavy deleted matches $M_{\mbox{\footnotesize old}}$. Each newly matched edge $m\in M_{\mbox{\footnotesize new}}$ can therefore, for accounting purposes, be associated with the match in $M_{\mbox{\footnotesize old}}$ that previously owned it, and we will use $p(m)$ to indicate this deleted match. Each matched edge $m\in M_{\mbox{\footnotesize new}}$ can contribute up to $r-1$ stolen deletes, one for each of its vertices except at least one incident on $p(m)$. Now each $e\in M_{\mbox{\footnotesize old}}$ can have at most $r$ newly matched edges incident on it (otherwise it would not be a matching). Therefore we can associate $r(r-1)$ stolen deletes with each $e\in M_{\mbox{\footnotesize old}}$. Furthermore we also associate the one bloated delete that might have caused each $e\in M_{\mbox{\footnotesize old}}$ to rise, and hence there are fewer than $r^{2}$ deletes associated with deleted edge $e$. This covers all the deletes.
We make the following observations:
  1. (1)
    $|E|=\sum_{e\in M_{\mbox{\footnotesize new}}}|S(e)|=\sum_{e\in M_{\mbox{\footnotesize old}}}|C(e)|$, since $|E|$ is the union of the cross edges in $M_{\mbox{\footnotesize old}}$ and all contribute to samples in $M_{\mbox{\footnotesize new}}$,
  2. (2)
    for $e\in M_{\mbox{\footnotesize old}}:|C(e)|\geq 4r^{2}2^{l(e)}$, since each is bloated or heavy, and
  3. (3)
    $S_{d}<r^{2}\sum_{e\in M_{\mbox{\footnotesize old}}}2^{l(e)+1}$, since each $e$ is associated with fewer than $r^{2}$ incident deletes and by Invariant 4.1.4 the level $l$ of each such delete is at most $l(e)$.
Together this gives:
$S_{a}\;\;=\;\;\sum_{e\in M_{\mbox{\footnotesize new}}}|S(e)|\;\;=\;\;\sum_{e\in M_{\mbox{\footnotesize old}}}|C(e)|\;\;\geq\;\;4r^{2}\sum_{e\in M_{\mbox{\footnotesize old}}}2^{l(e)}\;\;>\;\;2S_{d}$
We note that having the gap between levels be a constant factor ($\alpha=2$ in our case) instead of proportional to $r$ (as in Assadi and Solomon (Assadi and Solomon, 2021)) is important to the proof. In particular, in the last step of the proof of Lemma 5.6 we lose a factor of $\alpha$. Since we have $\alpha=2$, this brings the coefficient from $4$ to $2$ for us, which is fine. But if $\alpha$ was $\Theta(r)$, then the coefficient would be reduced from $4$ to $\Theta(4/r)$, which is too small to be useful. We could have compensated for $\alpha=\Theta(r)$ by increasing the cutoff for being heavy (in isHeavy) by a factor of $r$, to $\Theta(r^{3}2^{l(e)})$, but then the amortized cost of the overall algorithm would also have increased by a factor of $r$, to be $O(r^{4})$ per edge update. Thus, we save a factor of $r$ by using thinner levels. In AS, reducing the level gap from $\Theta(r)$ to $2$ is not necessary because they assign every new match to some specific deletes. This extra cost consideration in our algorithm comes from “shuffling” the sampled edges during the matching.
Lemma 5.7.
If we start and end with an empty graph, all work done by deleting matched edges can be covered by charging each natural deletion $O(r^{3}2^{l(e)})$ work in expectation.
Proof.
Let $S_{a}$ be the total sample size of all new epochs, $S_{i}$ the total sample size of all induced deletions and $S_{n}$ of natural deletes. From Lemma 5.6 we have that in each round of random settle the number of added samples is at least twice the number of samples from induced deletions. At the end of a batch update, the algorithm runs a insertEdges on the remaining unmatched edges. These are therefore not covered by the sample set of a random settle. Due to the termination condition, however, we know that the sample size of previously matched edges is twice as large as the number of the remaining edges. Let $|bloated|$ be the sample size of bloated remaining edges. Because these edges are bloated, they accrued a factor of at least $2r^{2}$ cross edges after rising. Thus, the sample size of deleted matches from the last round of random settle is at most $\frac{|\mbox{bloated}|}{2r^{2}}\leq\frac{|\mbox{bloated}|}{2}$. Therefore, the total sample space of induced deletions is at most the $\frac{1}{4}S_{a}$ from the final round plus the $\frac{1}{2}S_{a}$ from all other rounds, giving $S_{a}>\frac{4}{3}S_{i}$ within each batch deletion and hence also across all batch deletions. Since we start and end empty, we have that $S_{a}=S_{i}+S_{n}$. Putting these together we get $S_{n}>\frac{1}{3}S_{i}$.
Finally we note that the direct work we assign to each epoch is based on its level not its sample size. In particular, it is $cr^{3}2^{l(e)}$ for some constant $c$. Hence the cost per sample $C_{s}$ satisfies $\frac{1}{2}cr^{3}<C_{s}\leq cr^{3}$, since $2^{l(e)}\leq|S(e)|<2^{l(e)+1}$. This implies the per sample cost we assign to induced and natural epochs differs by at most a factor of two, which further implies that the cost assigned to natural deletions is at least a constant fraction of the cost assigned to induced deletions. The natural deletions can therefore cover the cost of induced deletions with a constant factor overhead. ∎
5.3. Overall Cost Bounds
We now consider the overall work and 0pt bounds. Firstly we note that if we have a bound for the work from a start where the graph is empty to an end when it is empty again, this implies a bound that is at least as good for any prefix at least half as long as the whole sequence. All runs are a prefix of at least half of an empty-to-empty sequence since we can just delete the edges and the number of final deletes is at most the number of prior operations. Hence we need only consider empty-to-empty runs to asymptotically bound the cost for the prefix.
Let $Q$ be the set of natural epochs. Let $d_{1},d_{2},\ldots,d_{T}$ be the set of deletions of sampled edges by the user. Note that different $d_{i}$’s may come from different instances of parallelGreedyMatch and that $T\leq N$. Suppose that we dynamically assign prices to sampled edges as described in Section 3.1.
Lemma 5.8.
The expected total natural sample space created is no more than the expected total price of all user deletes: $\mathbb{E}[\sum_{e\in Q}|S_{e}|]\leq\mathbb{E}[\sum_{t=1}^{T}\Phi^{\prime}(d_{t})]$.
Proof.
Because we start and end with an empty graph, note that the sum of the sample space created for all natural epochs is the same as the sum of the sample space of all natural epochs that is destroyed. Note that a natural epoch $e$ is destroyed by a user deletion and contributes $|S_{e}|$ to the total natural sample space destroyed.
Consider the user delete of an edge $e$. If $e$ is a cross edge, deleting it destroys no sample space, so we do not need to associate it to an epoch. If the edge is sampled, associate its delete with the epoch $e^{\prime}$ that owns it. Note that a delete is sampled iff it is early. Note that $e$ became sampled during some instance of parallelGreedyMatch. By Lemma 3.1, a sampled edge is owned by exactly one epoch within the instance of parallelGreedyMatch that sampled it. Because sampled edges do not participate in the algorithm until their matched edge is deleted, this edge will not join the sample space of any other epochs. Thus, the sample spaces of different instances of parallelGreedyMatch do not interfere with each other.
Let $e^{\prime}$ be a natural epoch. Because we end on an empty graph, $e^{\prime}$ is eventually deleted. Then applying Lemma 3.5, we get the total price of sampled deletes on $e^{\prime}$ is $|S_{e^{\prime}}|$. Since for each natural epoch, the sample space lost is no more than the price paid, the lemma follows. Note that the total price paid is not strictly equal to the total natural sample space lost because unmatched sampled deletes from induced epochs pay price but do not destroy natural sample space. ∎
Theorem 5.9.
For any sequence of batch updates starting and ending with an empty graph, and with $N$ edges added or deleted across all batches, the total work is $O(r^{3}N)$ in expectation.
Proof.
By Lemma 5.8, observe that $\mathbb{E}[\sum_{e\in Q}|S_{e}|]\leq\mathbb{E}[\sum_{t=1}^{T}\Phi^{\prime}(d_{t})]$. By Lemma 3.4, observe that $\mathbb{E}[\sum_{t=1}^{T}\Phi^{\prime}(d_{t})]\leq\sum_{t=1}^{T}\mathbb{E}[\Phi(d_{t})]\leq\sum_{t=1}^{T}2=2T$. By Lemma 5.7, the total work of the algorithm is bounded by $\sum_{e\in Q}Cr^{3}2^{l(e)}\leq\sum_{e\in Q}2Cr^{3}|S_{e}|\leq 2Cr^{3}(2T)\leq O(r^{3}N)$, where $C$ is the constant hidden by Lemma 5.7. ∎
Lemma 5.10.
The 0pt for every batch update is bounded by $O(\log^{3}m)$, where $m$ is the maximum number of edges at any point in the graph.
Proof.
The 0pt for randomSettle, insertEdges, and deleteMatchedEdges is $O(\log^{2}m)$. Specifically, the only steps in these functions that has 0pt more than $O(\log m)$ is the random greedy match, which has 0pt $O(\log^{2}m)$ by Lemma 3.3. The 0pt of deleteEdges is $O(\log^{2}m)$ up to but not including the while loop (dominated by the deleteMatchedEdges). The while loop will iterate at most $O(\log m)$ steps since $E^{\prime}$ cannot grow larger than $E$, and in each iteration $|E^{\prime}|$ is growing by a constant factor. Since the while loop makes a call to randomSettle, which has 0pt $O(\log^{2}m)$, the total 0pt is $O(\log^{3}m)$. ∎
6. Conclusion
In this work, we demonstrate that a maximal matching can be maintained in the parallel batch-dynamic setting against an oblivious adversary in $O(r^{3})$ expected work per update. We do so by using a black box random greedy maximal matching algorithm to handle unsettled edges at all levels at once, and arguing that we can charge the work in aggregate to natural epochs, which have small total weight.
An open question is whether the work can be improved to $O(r^{2})$. One key bottleneck in our work is the handling of stolen deletes. Since deleting a matched edge can cause $r^{2}$ stolen deletes, edges must have an $r^{2}$ factor before we permit them to random settle. Since edge updates naturally take $O(r)$ time for hyperedges, requiring an $r^{2}$ factor before random settling is cost prohibitive in the $O(r^{2})$ setting. It is open if it would be possible in a parallel algorithm to reduce the number of stolen deletes, or if an entirely different set of techniques would be needed.
Acknowledgements.  This work was supported by National Science Foundation grants CCF-2119352 and CCF-1919223.